# Anisotropic flow fluctuation as a possible signature of clustered nuclear geometry in O–O collisions at the Large Hadron Collider

### v1

@@ IV Summary
In this paper, we study the anisotropic flow coefficients in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using a hybrid of IPGlasma + MUSIC + iSS + UrQMD model and try to probe the final state effects of the presence of an exotic $\alpha$-cluster nuclear geometry in contrast to a uniform Woods-Saxon nuclear distribution inside a 16O nucleus. Here, we study the centrality dependence of anisotropic flow coefficients. In addition, for the first time, we report studies of elliptic flow fluctuations as a motivation to distinguish O–O collisions having $\alpha$-cluster geometry from a regular Woods-Saxon distribution. The observation of large triangular flow in the most central O–O collisions having an $\alpha$-clustered nuclei compared to a Woods-Saxon nuclear distribution is consistent with transport model predictions. In addition, we observe a significant centrality dependence of relative elliptic flow fluctuations in both the nuclear distributions. With this study, it is clear that the presence of an $\alpha$-cluster nuclear geometry has a significant effect on both elliptic and triangular flow. However, we note a distinctive feature appears when observables related to fluctuations are studied, i.e., triangular flow and elliptic flow fluctuations. While probing a nuclear density profile is a subject matter of low-energy nuclear physics, it will be very interesting to find an appropriate observable that can be measured in TeV nuclear collisions to do the job. This makes the present study more interesting.
@@ Acknowledgements
S. P. acknowledges the doctoral fellowship from UGC, Government of India. R. S. sincerely acknowledges the DAE-DST, Government of India funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021-IITI (E-37123). G. G. B. gratefully acknowledges the Hungarian National Research, Development and Innovation Office (NKFIH) under Contracts No. OTKA K135515, No. NKFIH NEMZ_KI-2022-00031 and Wigner Scientific Computing Laboratory (WSCLAB, the former Wigner GPU Laboratory). The authors gratefully acknowledge the MoU between IIT Indore and Wigner Research Centre for Physics (WRCP), Hungary, for the techno-scientific international cooperation.
@@
.1 Centrality versus impact parameter
@@
Figure 8: Average value of impact parameter as a function of centrality class in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD for a Woods-Saxon and $\alpha$-cluster nuclear density profile.
@@
Figure 8 shows the average value of the impact parameter ($\langle b\rangle$) as a function of centrality class in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD for a Woods-Saxon and $\alpha$-cluster nuclear density profiles. As expected, one finds a significant dependence of $\langle b\rangle$ on collision centrality selection based on the final state-charged particle multiplicity in the V0 acceptance region, i.e., the collisions having a lower impact parameter tend to produce a higher charged particle yield at the final state. In addition, one finds a significant dependence of $\langle b\rangle$ on the nuclear density profile of the colliding oxygen nucleus to produce a similar charge particle yield in the final state. For the $\alpha$-cluster case, one finds that $\langle b\rangle$ is significantly smaller compared to a Woods-Saxon density profile because of the compact nature of $\alpha$-cluster nuclear structure, the effects of which are enhanced in the peripheral collisions. In addition, the compact nature of $\alpha$-cluster nuclear density profile is also reflected in the higher yield in most central collisions, as shown in Fig. 4.
.2 Four particle cumulants
@@
The upper panel of Fig. 9 shows the centrality dependence of four-particle cumulants ($c_{n}\{4\}$) as a function of collision centrality in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The lower panel depicts the ratio of $c_{n}\{4\}$ from Woods-Saxon to a $\alpha$-cluster nuclear density profile. One can note that, $v_{n}\{4\}$ is related with $c_{n}\{4\}$ using Eq. (15), thus a negative value of $c_{n}\{4\}$ results in a physically acceptable value of $v_{n}\{4\}$. However, a positive value of $c_{n}\{4\}$ leads to an imaginary value of $v_{n}\{4\}$, which sometimes occur when the system is dominated with non-flow contributions [93]. In the upper panel of Fig. 9, one finds a significant difference of $c_{n}\{4\}$ between two nuclear density profiles. For an $\alpha$-cluster case, $c_{2}\{4\}$ starts with a small negative value and starts to decrease further with an increase in collision centrality, i.e., towards (10-20)centrality, attains a minimum and goes on increasing, until attains a positive value beyond (30-40)centrality class. In contrast, $c_{2}\{4\}$ for a Woods-Saxon nuclear density profile starts a minimum value for the most central case, i.e., (0-10)centrality class, and starts to rise, becomes positive beyond (30-40)centrality. Interestingly, for both the nuclear density profiles, $c_{2}\{4\}$ attains a positive value beyond a similar (30-40)centrality class, which naively implies the dominance of non-flow contribution for events with low final state multiplicity. The positive values of $c_{2}\{4\}$ on and beyond (30-40)centrality class restricts us from the calculation of four-particle elliptic flow and its fluctuations below (30-40)centrality class. On the other hand, $c_{3}\{4\}$ for both the nuclear density profiles fluctuate around 0, with a much smaller value as compared to corresponding $c_{2}\{4\}$. The positive values of $c_{3}\{4\}$ correspond to large non-flow azimuthal correlations.
.3 $p_{\rm T}$-dependence of anisotropic flow
@@
Figure 10 shows the transverse momentum dependence of elliptic and triangular flow calculated using the two-particle Q-cumulants method in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles in the upper panel. The bottom panel shows the ratio of $v_{n}\{2\}(p_{\rm T})$ the Woods-Saxon density profile to $\alpha$-cluster density profile. One finds both $v_{2}\{2\}(p_{\rm T})$ and $v_{3}\{2\}(p_{\rm T})$ rise linearly with an increase in $p_{\rm T}$, which is a exclusive feature of hydro-based models. In addition, one observes that for the most central (0-10)class, $v_{2}\{2\}(p_{\rm T})$ and $v_{3}\{2\}(p_{\rm T})$ are closer to one another, where the separation further increases as one moves towards the mid-central collisions. This is because the contributions for $v_{2}\{2\}(p_{\rm T})$ and $v_{3}\{2\}(p_{\rm T})$ are the consequences of the interplay between the geometry of the overlap region and the eccentricity fluctuations during the collisions, etc., where the eccentricity fluctuations dominate the contributions to the anisotropic flow in the central collisions. Another interesting observation is that, in the most central collisions, $v_{n}\{2\}(p_{\rm T})$ for the $\alpha$-cluster case is larger compared to corresponding values from a Woods-Saxon distribution. This observation is consistent with the Fig 5. However, towards the mid-central collisions, $v_{n}\{2\}(p_{\rm T})$ from Woods-Saxon nuclear distribution is larger compared to an $\alpha$-cluster nuclear density profile. For the central collisions, the lower ratio plot of Woods-Saxon to an $\alpha$-cluster nuclear density profile for both elliptic and triangular flow shows a flat ratio, almost independent of $p_{\rm T}$. However, it no longer holds towards the mid-central collisions, indicating a $p_{\rm T}$ dependence of medium effects for different nuclear profiles to the anisotropic flow coefficients.
@@
In this work, we used publicly available software for data generation. The software for IPGlasma, MUSIC, iSS and UrQMD can be accessed using the web links provided in Refs. [99, 100, 101, 102].

### v2

@@ IV Summary
In this paper, we study the anisotropic flow coefficients in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using a hybrid of IPGlasma + MUSIC + iSS + UrQMD model and try to probe the final state effects of the presence of an exotic $\alpha$-cluster nuclear geometry in contrast to a uniform Woods-Saxon nuclear distribution inside a 16O nucleus. Here, we study the multiplicity class dependence of anisotropic flow coefficients. In addition, for the first time, we report studies of elliptic flow fluctuations with a motivation to distinguish O–O collisions having $\alpha$-cluster geometry from a regular Woods-Saxon distribution. The observation of large triangular flow in the highest multiplicity O–O collisions having an $\alpha$-clustered nuclei compared to a Woods-Saxon nuclear distribution is consistent with transport model predictions. In addition, we observe a significant multiplicity class dependence of relative elliptic flow fluctuations in both the nuclear distributions. With this study, it is clear that the presence of an $\alpha$-cluster nuclear geometry has a significant effect on both elliptic and triangular flow. However, we note a distinctive feature appears when observables related to fluctuations are studied, i.e., triangular flow and elliptic flow fluctuations. While probing a nuclear density profile is a subject matter of low-energy nuclear physics, it will be very interesting to find an appropriate observable that can be measured in TeV nuclear collisions to do the job. This makes the present study more interesting.
@@ Acknowledgements
S. P. acknowledges the doctoral fellowship from UGC, Government of India. R. S. sincerely acknowledges the DAE-DST, Government of India funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021-IITI (E-37123). G. G. B. gratefully acknowledges the Hungarian National Research, Development and Innovation Office (NKFIH) under Contracts No. OTKA K135515, No. NKFIH NEMZ_KI-2022-00031, “2024-1.2.5-TET-2024-00022” and Wigner Scientific Computing Laboratory (WSCLAB, the former Wigner GPU Laboratory). The authors gratefully acknowledge the MoU between IIT Indore and Wigner Research Centre for Physics (WRCP), Hungary, for the techno-scientific international cooperation.
@@
.1 Multiplicity class versus impact parameter
@@
Figure 8: Average value of impact parameter as a function of multiplicity class in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD for a Woods-Saxon and $\alpha$-cluster nuclear density profile.
@@
Figure 8 shows the average value of the impact parameter ($\langle b\rangle$) as a function of multiplicity class in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD for Woods-Saxon and $\alpha$-cluster nuclear density profiles. As expected, one finds a significant dependence of $\langle b\rangle$ on event selection with the final state-charged particle multiplicity in the V0 acceptance region of the ALICE detector, i.e., the collisions having a lower impact parameter tend to produce a higher charged particle yield at the final state. In addition, one finds a significant dependence of $\langle b\rangle$ on the nuclear density profile of the colliding 16O nucleus to produce a similar charge particle yield in the final state. For the $\alpha$-cluster case, one finds that $\langle b\rangle$ is significantly smaller compared to a Woods-Saxon density profile because of the compact nature of $\alpha$-cluster nuclear structure, the effects of which are enhanced in the low multiplicity events. In addition, the compact nature of $\alpha$-cluster nuclear density profile is also reflected in the higher yield in 0–10% multiplicity class, as shown in Fig. 4.
.2 Eccentricity and triangularity
Figure 9: Upper panel shows the average and RMS values of eccentricity and triangularity as a function of multiplicity class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The lower panel shows the ratios of eccentricity and triangularity from Woods-Saxon to the $\alpha$-cluster nuclear density profile. In both the upper and lower panels, the vertical axes are zero-suppressed.
The upper panel of Fig. 9 shows the average eccentricity ($\langle\epsilon_{2}\rangle$), RMS eccentricity ($\sqrt{\langle\epsilon_{2}^{2}\rangle}$), average triangularity ($\langle\epsilon_{3}\rangle$), and RMS triangularity ($\sqrt{\langle\epsilon_{3}^{2}\rangle}$) as a function of multiplicity class in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma. Here, one finds that the average eccentricity for both the nuclear density profiles increases when one moves towards the lower multiplicity events, except for the $\alpha$-cluster case, which decreases slightly toward the 50–90% multiplicity class. The rate of this increment with a decrease in charge particle multiplicity is lower for the $\alpha$-cluster case as compared to the Woods-Saxon nuclear density profile. One finds that the values of $\langle\epsilon_{2}\rangle$ for the $\alpha$-cluster case are higher than the corresponding values of Woods-Saxon nuclear density profiles for 0–40% multiplicity classes; thereafter, Woods-Saxon has a higher value of $\langle\epsilon_{2}\rangle$. Further, one observes the average triangularity increases as one moves from the highest multiplicity region to the lower multiplicity region for the Woods-Saxon case. In contrast, $\langle\epsilon_{3}\rangle$ shows an interesting trend with the decrease in the charged particle multiplicity for the $\alpha$-cluster case. Both $\sqrt{\langle\epsilon_{2}^{2}\rangle}$ and $\sqrt{\langle\epsilon_{3}^{2}\rangle}$ retain a similar multiplicity class and nuclear density profile dependence observed for $\langle\epsilon_{2}\rangle$ and $\langle\epsilon_{3}\rangle$. It is important to note that average and RMS values of triangularity decrease by about 25% and 20%, respectively, when going from highest multiplicity to lowest multiplicity events of $\alpha$-cluster density profile.
The lower panel of Fig. 9 shows the ratios of $\langle\epsilon_{2}\rangle$, $\langle\epsilon_{3}\rangle$, $\sqrt{\langle\epsilon_{2}^{2}\rangle}$, and $\sqrt{\langle\epsilon_{3}^{2}\rangle}$ in O–O collisions with Woods-Saxon nuclear density profile to $\alpha$-cluster case. Here, for the highest multiplicity events, i.e., 0–10% multiplicity class, the Woods-Saxon nuclear density profile has smaller average and RMS values of $\epsilon_{2}$ and $\epsilon_{3}$ as compared to the corresponding values of $\alpha$-cluster case. Slowly, as the final state multiplicity decreases, the values of $\langle\epsilon_{2}\rangle$, $\langle\epsilon_{3}\rangle$, $\sqrt{\langle\epsilon_{2}^{2}\rangle}$, and $\sqrt{\langle\epsilon_{3}^{2}\rangle}$ gradually increase and become higher for the Woods-Saxon nuclear profile as compared to the corresponding values of $\alpha$-cluster case. Here, one finds that the ratio of average and RMS values of eccentricities and triangularities of the Woods-Saxon nuclear density profile to $\alpha$-cluster case are close to each other.
Figure 10: Upper panel shows the ratios $\langle\epsilon_{3}\rangle/\langle\epsilon_{2}\rangle$ and $\sqrt{\langle\epsilon_{3}^{2}\rangle}/\sqrt{\langle\epsilon_{2}^{2}\rangle}$ as a function of multiplicity class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The lower panel shows the ratio from Woods-Saxon to the $\alpha$-cluster nuclear density profile. In both the upper and lower panels, the vertical axes are zero-suppressed.
The upper panel of Fig. 10 shows the ratios $\langle\epsilon_{3}\rangle/\langle\epsilon_{2}\rangle$ and $\sqrt{\langle\epsilon_{3}^{2}\rangle}/\sqrt{\langle\epsilon_{2}^{2}\rangle}$ as a function of multiplicity class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. Here, one finds that, from 0–10% to 10–20% multiplicity class, the values of both $\langle\epsilon_{3}\rangle/\langle\epsilon_{2}\rangle$ and $\sqrt{\langle\epsilon_{3}^{2}\rangle}/\sqrt{\langle\epsilon_{2}^{2}\rangle}$ decrease rapidly for the $\alpha$-cluster case in contrast to Woods-Saxon nuclear density profile. Further, the values of $\langle\epsilon_{3}\rangle/\langle\epsilon_{2}\rangle$ and $\sqrt{\langle\epsilon_{3}^{2}\rangle}/\sqrt{\langle\epsilon_{2}^{2}\rangle}$ for the Woods-Saxon nuclear profile are higher as compared to the $\alpha$-cluster case except for the 0–10% multiplicity class, as can be seen in the lower panel of Fig. 10.
.3 Four particle cumulants
@@
The upper panel of Fig. 11 shows the multiplicity class dependence of four-particle cumulants ($c_{n}\{4\}$) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The lower panel depicts the ratio of $c_{n}\{4\}$ from Woods-Saxon to a $\alpha$-cluster nuclear density profile. One can note that, $v_{n}\{4\}$ is related with $c_{n}\{4\}$ using Eq. (15), thus a negative value of $c_{n}\{4\}$ results in a physically acceptable value of $v_{n}\{4\}$. However, a positive value of $c_{n}\{4\}$ leads to an imaginary value of $v_{n}\{4\}$, which sometimes occur when the system is dominated with non-flow contributions [98]. In the upper panel of Fig. 11, one finds a significant difference of $c_{n}\{4\}$ between two nuclear density profiles. For an $\alpha$-cluster case, $c_{2}\{4\}$ starts with a small negative value in the 0–10% multiplicity class and starts to decrease towards 10–20% multiplicity class, attains a minimum and goes on increasing until it attains a positive value beyond 30–40% multiplicity class. In contrast, $c_{2}\{4\}$ for a Woods-Saxon nuclear density profile starts at a minimum value for the most central case, i.e., 0–10% multiplicity class, and starts to rise, becomes positive beyond 30–40% multiplicity class. Interestingly, for both the nuclear density profiles, $c_{2}\{4\}$ attains a positive value beyond a similar 30–40% multiplicity class, which naively implies the dominance of non-flow contribution for events with low final state multiplicity. The positive values of $c_{2}\{4\}$ on and beyond 30–40% multiplicity class restrict us from the calculation of four-particle elliptic flow and its fluctuations below 30–40% multiplicity class. On the other hand, $c_{3}\{4\}$ for both the nuclear density profiles fluctuate around 0, with a much smaller value as compared to corresponding $c_{2}\{4\}$. The positive values of $c_{3}\{4\}$ correspond to large non-flow azimuthal correlations.
.4 $p_{\rm T}$-dependence of anisotropic flow
@@
Figure 12 shows the transverse momentum dependence of elliptic and triangular flow calculated using the two-particle Q-cumulants method in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles in the upper panel. The bottom panel shows the ratio of $v_{n}\{2\}(p_{\rm T})$ from the Woods-Saxon density profile to $\alpha$-cluster density profile. One finds both $v_{2}\{2,|\Delta\eta|>1.0\}(p_{\rm T})$ and $v_{3}\{2,|\Delta\eta|>1.0\}(p_{\rm T})$ rise linearly with an increase in $p_{\rm T}$, which is an exclusive feature of hydro-based models. In addition, one observes that for the 0–10% class, $v_{2}\{2,|\Delta\eta|>1.0\}(p_{\rm T})$ and $v_{3}\{2,|\Delta\eta|>1.0\}(p_{\rm T})$ are closer to one another, where the separation further increases as one moves towards 20–30% and 30–50% multiplicity classes. This is because the contributions for $v_{2}\{2,|\Delta\eta|>1.0\}(p_{\rm T})$ and $v_{3}\{2,|\Delta\eta|>1.0\}(p_{\rm T})$ are the consequences of the interplay between the geometry of the overlap region and the eccentricity fluctuations during the collisions, etc., where the eccentricity fluctuations dominate the contributions to the anisotropic flow in the central collisions. Another interesting observation is that in 0–10% multiplicity class, $v_{n}\{2\}(p_{\rm T})$ for the $\alpha$-cluster case is larger compared to corresponding values from a Woods-Saxon distribution. This observation is consistent with the Fig 5. However, for 20–30% and 30–50% multiplicity classes, $v_{n}\{2\}(p_{\rm T})$ from Woods-Saxon nuclear distribution is larger compared to an $\alpha$-cluster nuclear density profile. For the 0–10% multiplicity class, the lower ratio plot of Woods-Saxon to an $\alpha$-cluster nuclear density profile for both elliptic and triangular flow shows a flat ratio, almost independent of $p_{\rm T}$. However, it no longer holds towards the lower multiplicity classes, indicating a $p_{\rm T}$ dependence of medium effects for different nuclear profiles to the anisotropic flow coefficients.
@@
In this work, we used publicly available software for data generation. The software for IPGlasma, MUSIC, iSS and UrQMD can be accessed using the web links provided in Refs. [106, 107, 108, 109], respectively.